USING THE PARAMETER VARIATION METHOD,

Find the general solution of the differential equations taking into account the initial conditions.

Note: only determine all the matrices W in relation to the particular answer Yp without calculating them

yiv + 2y + y = 3t + 4 ; y(0) = y(0) = 0 et y(0) = y(0) = 1
Answer #1

(ن) my yov + 2y + y = 3+ ++ For C.F.; Put you + 2y+y=0 and find Cifi solution e complementary function). Auxiliary equation1/-tiv) Since Agaiм iSwiki,4 с. : un 4 + 1,42 +33 +uv yu 44 +u, 4, 4 и 93 + 4 - 0 ч ч + ч, 4,vy tu«Ч-о Че -- ц, ч, tv,Ч»» » » Из » Чч ам. доlutіоnа o 4 | 24 so Lyi - О - ) We left with; u + ч, ч, + , + ч уч Зt + 4 combining this equationDetermivaut of this «Не м «1. Ч. 2. » 03 , Чч is whanskian w of by cramens Full ча Ligh j f Wi Е: 3:(+9 Po W fo - | (-yh-i (

We are not supposed to solve it any further as per the note says that no need to calculate the matrices, so i am leaving the solution till here but without solving them we can't find the value of uj's and hence can't find particular integral and hence general solution. So to find the general solution first find out the determinants and then value of derivatives of uj's and by integrating find value of uj's and hence by putting them in yp find particular integral. And the general solution is y= yc+yp.